OpenStudy (littlenugget):

A power boat and a raft both left dock A on a river and headed downstream. The raft drifted at the speed of the river current. The power boat maintained a constant speed with respect to the river. The power boat reached dock B downriver, then immediately turned and traveled back upriver. It eventually met the raft on the river 9 hours after leaving dock A. How many hours did it take the power boat to go from A to B?

OpenStudy (anonymous):

let "U" be the speed of the river current let "v" be the speed of the speedboat let "L" be the total distance between A and B let "x" be the total distance from point A to the raft

OpenStudy (anonymous):

making the assumption that the raft has not reached point B in 9 hours

OpenStudy (anonymous):



OpenStudy (anonymous):

in the 9 hours, the raft only travelled x distance meaning you can determine speed of the raft/ speed of the river current

OpenStudy (anonymous):

in the 9 hours, the speed boat travels from point A - B, then travels from B to the location of the raft this means the total distance travelled by the speedboat would be L + (L-x)

OpenStudy (anonymous):

the speedboat speed when travelling from point A to point B would be U+V and the speedboat's speed when travelling upriver would be V-U "making the assumption that the speed of the speedboat is faster than the speed of the current of the river"

OpenStudy (anonymous):

let "t1" be the amount of time taken for the speedboat to go from point A to point B using the knowledge given about the speedboat, you can create 2 equations L= t1 (U+v) and L-x= (9-t1)( V-U) the first equation is the distance speed equation for the speedboat as i ts going from point A - point B the 2nd equation is the distance speed equation for the speedboat going from point B to the position of the raft at this point, i think its just a bunch of substitution and solving for t1